OpenStudy (anonymous):

Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is three times its width; if the area of the garden is 75 square meters, what are its dimensions?

OpenStudy (anonymous):

3w*w=75 w^2=25 w=5 l=5*3=15

OpenStudy (angela210793):